OpenStudy (anonymous):

Each side of a square loop of wire measures 2.0 cm. A magnetic field of 0.044 T perpendicular to the loop changes to zero in 0.10 s. What average emf is induced in the coil during this change? (Points : 3) 1.8 V 0.088 V 0.88 V 0.00018 V I ttink its D

OpenStudy (anonymous):

@SolomonZelman @aaronq @RedNeckOutLaw @confluxepic

OpenStudy (irishboy123):

Faraday's Law: \[\mathcal {E} = - \frac{d \phi}{dt} = - \frac{d \ (BA)}{dt} = - A \ \frac{dB}{dt}\] dB/dt = -0.044/0.1 A = (0.02)^2

OpenStudy (shamim):

U know magnetic flux=B.A=BA*cos 0=BA*1=BA

OpenStudy (shamim):

Do u know the angle between B nd A